MAGNETIC CIRCULAR DICHROISM STUDY OF SPIN-ORBIT SPLITTING IN ALKALI DIMERS IN THE PRESENCE OF A HELIUM SURFACE

GERALD AUBÖCK, JOHANN NAGL, \underline{CARLO CALLEGARI}, and WOLFGANG E. ERNST, Institute of Experimental Physics, TU Graz, Petersgasse 16, 8010 Graz, Austria/EU.

We present the application of Magnetic Circular Dichroism (MCD) to the determination of the level structure of orbitallydegenerate electronic levels of molecules under the perturbation of a He droplet, which had so far escaped assignment. Our target system is the (1)^{3}Π_{g} [(2)^{3}Π if heteronuclear] manifold of alkali dimers.^{a} The perturbation due to the droplet breaks the rotational symmetry around the internuclear axis of the diatom and causes a splitting of this orbitally-degenerate manifold. We extend to He droplets a model previously used to interpret the matrix spectra of the NH radical.^{b} With a small number of physically reasonable parameters the model accounts for the essential features of laser-induced fluorescence (LIF) and magnetic circular dichroism (MCD) spectra of Rb_{2} and K_{2}. MCD spectra are not only essential to the correct assignment of the observed structure, but also allow a determination of the populations of Zeeman sublevels in the ground state and thus a measurement of the surface temperature of the droplet. The latter agrees with the accepted temperature, 0.37 K, measured in the interior of a droplet.

aG. Auböck, J. Nagl, C. Callegari, and W. E. Ernst, J. Phys. Chem. A 111, 7404 (2007).

bV. S. Langford and B. E. Williamson J. Phys. Chem. A 2415, 102 (1998).